OpenStudy (anonymous):

In the function, How much of the substance will be present after 20 years? Fuction is attatched below

OpenStudy (anonymous):

what function?

OpenStudy (anonymous):

@myko I posted it

OpenStudy (anonymous):

\[A(20)=4(1/2)^{1/5*20}=2^{4}=2*2*2*2=16\]

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

yw